OpenStudy (anonymous):

equation of the tangent line to the curve y=e^(1-x)+3/(x^2)+1

OpenStudy (anonymous):

at which point?

OpenStudy (anonymous):

Stfu b!tch

OpenStudy (anonymous):

@Mathmuse closed intereval (0,3)

OpenStudy (anonymous):

the equation will change based on which point on the interval you are evaluating because the slope at each point (the derivative) will be different

OpenStudy (anonymous):

Do you even lift 

OpenStudy (anonymous):

? \[\large y= e^{(1-x)}+\frac{3}{x^2+1}\]

OpenStudy (anonymous):

the e^(1-x) is on the line with the +3

OpenStudy (anonymous):

@Mathmuse

OpenStudy (anonymous):

stfu

OpenStudy (anonymous):

\[\large y= \frac{e^{(1-x)}+3}{x^2+1}\]

OpenStudy (anonymous):

yes, need the equation of the tangent line at (1,2)

OpenStudy (anonymous):

ahhh..ok (1,2) you'll need to fill in the following form \[\large y - y_1 = m(x-x_1)\] where the values with subscripts can be those two points (ie x1 = 1, y1 = 2) and the m is the slope. this is found by taking the derivative of the function and subbing those same two x,y values

OpenStudy (anonymous):

@ssaid2 clear as mud?